OpenStudy (anonymous):

Two end points of a line segment are (6, 8) and (-4, -2). What are the coordinates of the point on the line through which its bisector passes? (2, 6) (1, 3) (7, -3) (2, 2)

OpenStudy (anonymous):

First you should get the equation for the original line. Then you should be able to figure out the equation of the bisecting line. Then just see which of the answers fits into that equation

OpenStudy (anonymous):

\[Midpoint =\left( \frac{ x1+x2 }{ 2 } , \frac{ y1+y2 }{ 2 }\right)\] Therefore: \[\left( \frac{ -4+6 }{ 2 } , \frac{ -2+8 }{ 2 } \right) = \left( \frac{ 2 }{ 2 } , \frac{ 6 }{ 2 }\right)\] So the bisection line at the midpoint would be at point (1,3)

hartnn (hartnn):

the bisector will pass through mid-point of those 2 points. the formula for co-ordinates for mid-point(m,n) of points (x1,y1)and (x2,y2) is m=(x1+x2)/2 n=(y1+y2)/2 just plug in x1=6,y1=8,x2=-4,y2=-24 to get the midpoint.